### v3

# From Flatland to Space: Teaching Vision-Language Models to Perceive and Reason in 3D

Abstract
Recent advances in LVLMs have improved vision-language understanding, but they still struggle with spatial perception, limiting their ability to reason about complex 3D scenes. Unlike previous approaches that incorporate 3D representations into models to improve spatial understanding, we aim to unlock the potential of VLMs by leveraging spatially relevant image data. To this end, we introduce a novel 2D spatial data generation and annotation pipeline built upon scene data with 3D ground-truth. This pipeline enables the creation of a diverse set of spatial tasks, ranging from basic perception tasks to more complex reasoning tasks. Leveraging this pipeline, we construct SPAR-7M, a large-scale dataset generated from thousands of scenes across multiple public datasets. In addition, we introduce SPAR-Bench, a benchmark designed to offer a more comprehensive evaluation of spatial capabilities compared to existing spatial benchmarks, supporting both single-view and multi-view inputs. Training on both SPAR-7M and large-scale 2D datasets enables our models to achieve state-of-the-art performance on 2D spatial benchmarks. Further fine-tuning on 3D task-specific datasets yields competitive results, underscoring the effectiveness of our dataset in enhancing spatial reasoning.
Figure 1: Overview of our Spatial Perception And Reasoning (SPAR) dataset and benchmark. Our dataset is sourced from 4,500 scenes and comprises 33 spatial tasks spanning single-view, multi-view, and video settings. Our benchmark includes over 7,000 carefully curated high-quality samples to comprehensively evaluate the spatial understanding capabilities of existing models.
1 Introduction
Dataset Perception Reasoning Single view Multi $\dagger$ Video Task types QA quantity
ScanQA [4] - 27k
SQA3D [42] 16 33k
3D-LLM [25] 9 659k
LEO [29] 9 513k
SceneVerse [31] 3 2.5M
MMScan [40] 20 6.9M
SPAR-7M 33 7M
Table 1: Comparison of spatial understanding datasets. $\dagger$ Multi-view questions involve 3–5 images per QA pair.
Recent advances in Large Vision-Language Models (LVLMs) have significantly enhanced vision-language understanding [3, 21, 35, 34, 5]. However, to empower VLMs to better interact with the physical world, spatial understanding is essential, serving as a cornerstone of embodied intelligence. Nevertheless, current VLMs exhibit limited spatial perception, often struggling with comprehensive understanding and reasoning about spatial relationships in complex scenes [10, 19, 25, 21, 14, 56].
A recent line of research explores enhancing spatial understanding by incorporating 3D representations of scenes [25, 14, 56, 28, 23]. However, these methods encounter several challenges. First, 3D data are scarce, with suboptimal quality and uneven distribution, making them unsuitable for large-scale pretraining. Second, designing a model (e.g., a point cloud encoder) to effectively capture 3D information is inherently challenging, particularly since complete 3D representations are often inaccessible in real-world scenarios. Moreover, these approaches modify the network architecture, hindering seamless integration with existing VLM frameworks [5, 35].
This raises the question: Is it truly necessary to incorporate 3D representations into VLMs? For humans, spatial understanding of the relationship between surrounding objects is innate, relying on the ability to observe and implicitly reconstruct the entire space from 2D observations. Motivated by this capability, we seek to empower VLMs to acquire spatial understanding directly from large-scale 2D data (e.g., image inputs), enabling them to perform spatial reasoning without explicit 3D supervision. To facilitate this, we construct a spatially relevant dataset tailored for training and evaluation. Additionally, to enable VLMs to tackle a broader range of 3D tasks, we design a multi-view-based 3D grounding approach.
We categorize spatial understanding into two levels: spatial perception, which encompasses fundamental spatial concepts such as depth, distance, relative positions, and camera viewpoints, representing low-level cognitive abilities; and spatial reasoning, which involves higher-order spatial understanding, requiring the ability to infer and construct 3D scenes implicitly. A critical aspect of this advanced ability is view matching, particularly in multi-view contexts.
Building on these two levels of spatial understanding, we introduce SPAR-7M, a dataset encompassing 33 tasks that cover a diverse range of complexities. These include basic perception tasks such as depth and distance estimation, more advanced reasoning tasks like cross-view object matching, and high-level spatial cognition tasks such as spatial imagination. To systematically organize these tasks, we classify them into three levels: low, medium, and high, reflecting a progression from Spatial Perception to Spatial Reasoning. To generate the necessary training data, we introduce a 2D data generation and annotation pipeline that derives rich supervision from 3D scene data. The resulting dataset, SPAR-7M, provides precise 3D ground-truth annotations, significantly improving both task diversity and reliability. Unlike existing approaches that rely solely on 2D annotations [10, 19], our dataset enables a more structured and accurate assessment of spatial understanding. Tab. 1 illustrates the diversity of input formats and task types within our dataset.
Despite the growing body of spatial benchmarks, many focus predominantly on high-level spatial reasoning, often overlooking fundamental spatial perception [24, 41, 40]. Furthermore, some benchmarks, particularly those using internet-collected images, struggle to effectively evaluate multi-view tasks, while others based on video inputs are not well-aligned with typical model inputs [52]. To address these limitations, we select 20 tasks from the SPAR-7M validation set. After carefully manual validation and filtering, we created SPAR-Bench, a comprehensive benchmark consisting of 7,207 questions that span tasks from basic perception to complex reasoning. SPAR-Bench supports both single-view and multi-view inputs, offering a more holistic and diverse evaluation of spatial understanding.
Our main contributions are as follows: (i) We introduce a 2D spatial QA generation pipeline leveraging 3D ground truth, enabling the creation of both single-view and multi-view data. Additionally, we present SPAR-7M, a comprehensive dataset covering a wide range of tasks from spatial perception to reasoning. (ii) We present SPAR-Bench, a benchmark for evaluating the spatial capabilities of existing models. (iii) By training models on both SPAR-7M and large-scale 2D internet datasets, we achieve state-of-the-art performance across multiple 2D spatial benchmarks. Moreover, even with 2D-only inputs, further fine-tuning on 3D task-specific datasets yields competitive results. We will open-source SPAR-7M and SPAR-Bench, facilitating further research in spatial understanding.
2 Related work
Datasets for spatial understanding.
Existing multimodal datasets primarily provide single-view or limited multi-view images, often designed for image-text retrieval or basic visual question answering, making them insufficient for advanced 3D reasoning [33, 54, 39, 38]. Some datasets attempt to address this by collecting large-scale images from the internet, but they lack precise 3D annotations, relying instead on estimation models, which limits their effectiveness in multi-view and spatial reasoning tasks [10, 19]. Others incorporate image sequences or 3D scene representations for object counting, spatial relationship reasoning, or grounding, but they mainly focus on high-level semantics while neglecting low-level geometric cues such as depth, distance, and camera parameters [4, 42, 16]. Recent efforts [47] explore spatial reasoning from a robotics perspective, emphasizing reference-frame understanding and grounded spatial relations in real-world scenes. While highly relevant to embodied agents, these datasets focus on a narrower set of reasoning types and lack the task diversity and cognitive stratification we aim for. To overcome these limitations, we propose a 3D-ground-truth-based data generation pipeline capable of rendering single-view, multi-view, and video-based data. Our approach ensures precise spatial supervision, covering a broad spectrum of tasks from basic perception to advanced reasoning, structured into three difficulty levels. This provides a more comprehensive and scalable solution compared to existing datasets.
Benchmarks for evaluating spatial understanding.
Existing spatial understanding benchmarks primarily focus on high-level reasoning while often neglecting fundamental spatial perception tasks [41, 24, 40, 4, 42, 52]. Most of them rely on either single-view images or video inputs, with limited consideration of multi-view settings, despite their ability to better represent local 3D structures. Single-view images restrict the complexity of questions that can be formulated and provide an incomplete assessment of VLMs’ spatial reasoning capabilities. While some benchmarks use video inputs [52], many VLMs are not optimized for video processing, but this does not necessarily mean they lack spatial reasoning abilities. To address these limitations, we introduce SPAR-Bench, a more comprehensive benchmark that encompasses both fundamental and advanced spatial tasks. It supports both single-view and multi-view inputs, enabling a more systematic and reliable evaluation of spatial understanding.
3 Method
Our research aims to equip VLMs with robust and accurate spatial understanding, focusing on both fundamental 3D perception and higher-order spatial reasoning and imagination. To address the challenges posed by limited data, we propose a novel data generation pipeline that creates a diverse range of spatial understanding tasks. Our pipeline leverages 3D scenes from ScanNet [20], ScanNet++ [53], and Structured3D [55], ensuring a broad and diverse representation of indoor environments. Along with this, we introduce a new benchmark to provide a comprehensive evaluation of models’ spatial capabilities. In the following sections, we first detail our data generation pipeline (Sec. 3.1) and the key characteristics of SPAR-7M (Sec. 3.2). We then introduce SPAR-Bench (Sec. 3.3), followed by an analysis of data quality and the manual filtering process used to ensure benchmark reliability (Sec. 3.4). Additionally, in Sec. 3.5, we present a multi-view-based 3D grounding approach designed to enhance VLMs’ ability to tackle a broader range of 3D tasks.
3.1 3D-driven data generation pipeline
This section introduces our proposed data generation pipeline. The input consists of the scene mesh, 3D bounding boxes of objects, the intrinsic and extrinsic of camera sequences, and the image sequences. Through a series of steps, including filtering, extracting 3D information, and task construction, we generate diverse spatial understanding data. The key steps of this process are detailed below, with the overall workflow illustrated in Fig. 2.
Figure 2: Data generation pipeline. Our pipeline consists of three main parts: (1) Filtering redundant images by sampling approach. (2) Extracting scene 3D information and creating a unified data structure to support multi-view cross-retrieval and annotation. (3) Leveraging metadata to automatically generate a variety of tasks and forms of QA data.
Filtering and subsampling.
In the first step, we focus on reducing data redundancy by filtering out images with minimal camera movement. Specifically, each image is associated with a camera pose (position and orientation). If an image’s camera position is within a distance threshold of a retained image and its orientation differs by less than an angular threshold, it is considered redundant and removed. The whole process is similar to keyframing and the implementation details can be found in the appendix  6.2.
Extracting 3D information.
In the second step, we extract global 3D information from the scene and create a unified data structure to support multi-view cross-retrieval and annotation. This structure facilitates dataset alignment and the generation of downstream tasks. Each scene is represented by two types of entities: Image records and object records.
  • Image record stores metadata for each image, such as its index, camera intrinsics and extrinsics, image dimensions, and the list of objects present. Each object is linked to its 3D bounding box, 3D center point in world coordinates, and its 2D bounding box on the image plane.
  • Object record contains global information for all objects in the scene, including their ID, category, and a list of images where they appear (referencing the corresponding image indices).
This structure creates a unified mapping between image and object data in world coordinates, simplifying the alignment and management of data from different sources. It also lays the foundation for multi-dataset annotations and task generation. Additional technical details can be found in the appendix 6.3.
Multi-tasks generation.
In the third step, we use the complete 3D scene information to automatically generate a variety of tasks and corresponding QA pairs. Each task is configured with key details such as task type (e.g., distance prediction, spatial relations), input image count, question format, and the object types involved (object-to-object or object-to-observer relations).
For each task, we follow a specific pipeline to generate annotations, ensuring flexibility to extend or customize tasks as needed. After generating the annotations, placeholders in the template (e.g., [label_A], [label_B], [left_right]) are replaced with actual object labels and computed spatial information. The templates are initially created by humans, expanded with the help of GPT, and then refined through manual selection to ensure diversity and accuracy.
While the process for generating tasks is consistent, each of the 31 tasks follows its own unique procedure based on its specific requirements. Detailed information about task pipelines, QA templates, and GPT prompting can be found in the appendix 6.4.
Figure 3: Dataset and benchmark visualization. We visualize some samples from our proposed SPAR-7M and SPAR-Bench. We show the diversity and richness of our dataset, which covers different task types, input types, and object types.
3.2 Spatial dataset construction
Based on our data generation pipeline (Sec. 3.1), we construct SPAR-7M, a large-scale dataset designed for spatial understanding. SPAR-7M consists of approximately 4K indoor 3D scenes from diverse sources, covering 33 task types with over 7 million QA pairs. Each QA pair is paired with camera intrinsics, extrinsics, and depth information, enabling both multi-view reasoning and 3D spatial inference. The dataset includes tasks ranging from basic spatial perception (e.g., depth estimation, distance measurement) to high-level spatial reasoning (e.g., camera position prediction, multi-view spatial relations) (Fig. 4).
To comprehensively evaluate different aspects of spatial understanding, we incorporate three types of input formats: single-view, multi-view, and video. Single-view tasks test a model’s ability to infer spatial properties from a single image, while multi-view tasks require integrating spatial information across different viewpoints. The video-based setting goes beyond individual frames to introduce higher-level spatial perception and scene-level reasoning, challenging the model to understand the spatial structure of a more complex environment. While not all tasks include every input format, this diversity ensures a broad assessment of spatial reasoning across various contexts.
In addition to varying input formats, we categorize tasks based on their spatial relationships into object-object (OO) and object-camera (OC) types. OO tasks focus on the relative positioning of objects, such as determining whether one object is to the left, right, or behind another. OC tasks, on the other hand, emphasize the relationship between objects and the camera, requiring models to infer viewpoint-dependent properties. For multi-image inputs, one image is randomly selected as the primary viewpoint, serving as the reference for spatial reasoning.
3.3 Spatial benchmark construction
To systematically evaluate spatial understanding, we introduce SPAR-Bench, a benchmark designed to assess models across diverse spatial reasoning tasks and input modalities. Unlike existing benchmarks that primarily focus on high-level reasoning or rely on video-based inputs, SPAR-Bench provides a more balanced evaluation by covering both spatial perception and reasoning in single-view and multi-view settings.
SPAR-Bench consists of 20 tasks covering depth estimation, distance prediction, spatial relation reasoning, and spatial imagination. These tasks are sampled from the SPAR-7M test set, with 400 QA pairs per task. After manual quality checking, the final benchmark includes 7207 QA pairs. Each task is formulated as either numerical prediction (fill-in-the-blank) or classification (multiple-choice selection), evaluated using Mean Relative Absolute Error (MRA) and accuracy, respectively.
To capture different levels of spatial understanding, single-view tasks focus on spatial inference from a single image, while multi-view tasks require reasoning across multiple perspectives, such as object matching and camera pose inference. A full list of tasks and details on data construction are provided in the appendix 6.1. We visualize the dataset and benchmark in  Fig. 3. More detailed QA of datasets can be found in appendix 6.5.
Figure 4: Dataset statistics. Our Dataset can be categorized into high, medium, and low three levels, totaling 33 subtasks.
3.4 Data quality check
To ensure the reliability of our proposed benchmark, we conduct a human verification process. Specifically, we sample 400 questions per task in the validation set, forming an 8,000-question benchmark. These questions undergo manual inspection, taking approximately 140 human hours. After filtering out problematic cases, 7,207 questions were retained, resulting in a dataset acceptance rate of 90.1%.
During verification, we focus on identifying vague question stems, incorrect or misleading choices, ambiguous answers, and factual inconsistencies. Errors are marked, discussed, and either corrected or removed. This rigorous process ensures that our benchmark maintains high quality and reliability.
Figure 5: Our 3D grounding pipeline via VLM. For multi-view inputs, we decompose the task into frame selection and mono-3D grounding on the selected frame.
3.5 3D grounding with VLMs
Expressing an object’s spatial position within a scene is a crucial capability. Previous methods primarily relied on point cloud regression or classifying proposal bounding boxes. For video inputs, temproal grounding in VLM [26, 44] is widely explored, [51] use VLM as a frame selector to achieve 3D grounding with a series of tools. In contrast, we propose a novel 3D grounding approach that reformulates this task as a text generation problem, enabling seamless integration with VLMs. Specifically, we transform the 3D grounding task into a frame selection followed by mono-3D grounding. In the case of multi-view input, an object can be localized within a visible frame, allowing us to infer its UV coordinates, depth, and size, achieving accurate 3D localization through the camera pose of that frame. While this method already enables the localization of a 3D bounding box, we further introduce an optional refinement process. By optimally matching the predicted 3D bounding box with the proposal bounding box, we obtain the final refined result. The overall grounding pipeline is illustrated in Fig. 5.
4 Experiments
VSI-Bench [52] CV-Bench   [49] BLINK [24] 3DSRBench [41] Seed-Image [33] MME [22] MMBench [38] RealWorldQA [43] TextVQA [46]
Methods 2D 3D
GPT-4v [2] - 64.3 73.8 51.1 - 71.6 1927 75.0 61.4 77.4
GPT-4o [2] 34.0 - - 60.0 45.3 77.1 2310 83.4 75.4 -
Cambrain-8B [49] - 72.3 72.0 - - 74.7 1547 75.9 64.2 77.8
LLaVA-OV-7B [34] 25.3 - - 48.2 44.1 - 1998 80.8 66.3 -
InternVL2-8B [18] 34.6 - - - - 75.4 2215 81.7 - 77.4
Base VLM 32.4 74.2 78.5 46.6 58.3 76.5 2323 84.5 66.7 68.7
+EMOVA-2M 24.5 66.3 64.8 42.4 55.3 73.8 2186 80.2 63.1 63.8
+SPAR-mix 41.1 72.3 89.1 43.9 57.5 73.2 2163 79.9 64.7 62.9
Table 2: Comparison of model performance on various spatial understanding and 2D gerneral benchmarks. The base VLM is InternVL2.5-8B, and we use dynamic patch 3 during training and inference.
QA source Hypersim [45] SUNRGBD [48] nuScenes [8]
Base VLM 73 87 62
+SPAR-mix 93 96 80
Table 3: Performance improvement on out-of-domain datasets in CV-Bench 3D.
In this section, we assess the effectiveness of our dataset and investigate the potential upper limits of spatial reasoning in VLMs from three key perspectives:
  • Performance on existing spatial benchmarks: We examine how pretraining on SPAR-7M improves model performance on image-based spatial benchmarks.
  • Evaluation on our in-domain benchmark SPAR-Bench: We assess the pretrained model’s capabilities using our proposed benchmark, SPAR-Bench, which provides a comprehensive evaluation of spatial understanding across various tasks.
  • Fine-tuning on 3D-specific tasks: To investigate the generalization ability of our approach, we perform supervised fine-tuning (SFT) on 3D-specific benchmarks, including spatial QA and 3D grounding tasks.
Furthermore, in Sec. 4.4, we analyze the impact of different data compositions and training strategies on model performance, offering insights into optimal configurations for enhancing spatial reasoning capabilities.
Pretraining.
To train a VLM with strong spatial capabilities while preserving generalization, we use a mixed pretraining dataset SPAR-mix. 40% of the data is uniformly sampled from our dataset, SPAR-7M, while the remaining 60% is uniformly sampled from general 2D data, sourced from the EMOVA configuration. This general data includes tasks such as image-text QA, chart understanding, and OCR. We pretrain the model on a total of 2 million samples, using the base model InternVL2.5-8B.
4.1 Evaluation on existing 2D spatial benchmarks
CV-Bench [49].
The Cambrian Vision-Centric Benchmark (CV-Bench) is a large-scale dataset with 2,638 manually verified examples, designed to assess 2D spatial relationships and object counting. It repurposes standard vision datasets, formulating natural language questions to evaluate spatial reasoning. Visual prompts, such as bounding boxes, are used to reduce ambiguity and ensure reliable assessments.
VSI-Bench [52].
The Visual-Spatial Intelligence Benchmark (VSI-Bench) evaluates spatial reasoning from egocentric video data, containing 5,000+ QA pairs from 288 real-world videos sourced from ScanNet [20], ScanNet++ [53], and ARKitScenes [7]. It covers diverse indoor environments and leverages object-level annotations for question generation. VSI-Bench includes eight tasks across three categories: configurational reasoning (object counting, relative position, route planning), measurement estimation (object and room size, absolute distance), and spatiotemporal reasoning (appearance order in videos). All questions undergo manual review to ensure accuracy.
To evaluate the effectiveness of our dataset, we compare models pretrained on our SPAR-mix (a combination of our data and the EMOVA dataset) against those trained solely on 2M EMOVA dataset [12]. As shown in Tab. 2, Pretraining with our mixed dataset results in significant performance gains and achieves SOTA performance on both the CV-Bench 3D and VSI-Bench benchmarks, which are specifically designed to evaluate spatial reasoning capabilities.Moreover, compared to models trained exclusively on 2M EMOVA dataset, our approach maintains competitive 2D generalization performance, demonstrating that integrating our dataset effectively enhances 3D spatial reasoning while preserving 2D capabilities.
Out-of-domain generalization ability.
We also find that pretraining with our dataset enhances the model’s spatial understanding across diverse, non-homogeneous datasets. Although our dataset is primarily derived from indoor scene datasets such as ScanNet [20], ScanNet++ [53], and Structured3D [55], the pretrained model shows consistent improvements across all tasks constructed with datasets different from ours in CV-Bench 3D, as shown in Tab. 3. Notably, on NuScenes, an outdoor dataset, our approach yields an 18-point performance gain, demonstrating its strong generalization ability beyond indoor environments.
Methods
Rank
Avg.
Low
Depth-OC
Depth-OC-MV
Depth-OO
Depth-OO-MV
Dist-OC
Dist-OC-MV
Dist-OO
Dist-OO-MV
Medium
PosMatch
CamMotion
ViewChgI
High
DistI-OO
DistI-OO-MV
ObjRel-OC-MV
ObjRel-OO
ObjRel-OO-MV
SpImag-OC
SpImag-OC-MV
SpImag-OO
SpImag-OO-MV
Baseline
Chance Level (Random) - - - - - - - - - - - - 22.65 24.50 - 25.09 23.82 22.02 31.25 25.27 22.16 25.81 24.42 24.17 26.89
Chance Level (Frequency) - 32.74 31.19 43.09 43.51 17.38 13.05 41.90 30.99 27.40 32.17 38.25 29.01 26.75 59.00 32.29 52.94 50.60 28.25 26.92 26.59 26.34 26.74 26.49 25.77
SPAR-Bench (tiny) API
Human Level 1 67.27 55.31 72.75 74.25 28.75 36.25 78.25 52.25 66.5 33.50 72.32 92 64 60.97 76.22 80 94 70 92 80 78 82 50 60
GPT-4o [2] 3 36.39 29.25 53.80 45.00 15.00 13.60 37.40 34.40 23.40 24.40 24.93 30 16 28.80 45.11 64 64 58 46 46 32 44 30 22
Claude-3.7-Sonnet [1] 5 21.77 25.43 41.00 45.40 11.20 12.20 42.60 19.60 26.00 5.40 7.33 16 6 0.00 23.33 40 48 22 36 14 12 20 6 12
Qwen2-VL-72B [50] 4 35.62 35.28 45.40 49.80 13.80 10.00 54.60 49.40 36.80 22.40 23.39 42 18 10.16 40.00 60 68 50 38 44 18 28 18 36
Qwen2.5-VL-72B [6] 2 39.40 35.35 53.20 46.80 17.80 29.00 49.60 57.40 14.40 14.60 23.05 40 16 13.16 48.44 74 74 60 56 50 20 34 24 44
SPAR-Bench (full)
InternVL2-2B [18] 10 28.06 21.74 18.06 24.81 23.20 20.97 19.47 19.95 26.83 20.61 22.83 39.69 23.00 5.81 35.42 51.18 55.95 46.00 31.59 23.82 36.02 34.30 17.55 22.41
InternVL2-4B [18] 4 32.01 28.94 23.94 27.22 20.00 18.12 42.57 40.16 31.29 28.18 29.16 49.87 21.00 16.62 35.70 56.76 55.36 40.25 36.81 25.21 28.76 32.27 21.19 24.65
InternVL2-8B [18] 3 33.02 26.83 25.75 30.88 20.67 20.78 39.03 36.19 19.15 22.19 36.49 63.36 28.00 18.11 37.37 64.71 54.46 42.75 37.36 26.32 34.14 31.10 20.86 24.65
InternVL2.5-2B [17] 8 30.14 25.79 39.67 39.72 12.12 15.03 30.94 29.59 20.22 19.02 22.93 37.91 24.25 6.64 36.41 51.47 56.85 50.25 33.79 24.10 27.15 35.17 26.49 22.41
InternVL2.5-4B [17] 7 30.55 25.66 29.06 32.97 21.77 16.83 20.84 26.85 28.13 28.79 29.75 47.07 33.25 8.92 35.16 54.12 58.93 35.50 29.67 34.63 24.73 31.39 19.21 28.29
InternVL2.5-8B [17] 1 36.28 29.46 25.78 29.31 23.79 18.76 46.82 42.68 22.62 25.89 31.88 61.32 28.00 6.32 43.80 59.71 56.85 51.75 44.23 41.55 36.56 41.57 22.52 39.50
LLaVA-OV-0.5B [34] 9 29.48 30.14 49.22 42.72 18.04 14.92 31.48 25.67 28.98 30.10 15.89 24.43 21.75 1.50 33.42 50.88 50.00 32.00 27.75 26.04 30.91 34.01 24.50 24.65
LLaVA-OV-7B [34] 5 31.20 21.79 30.33 26.94 18.58 13.87 10.43 13.64 31.24 29.29 26.13 38.68 30.25 9.47 40.14 56.47 55.06 37.25 48.63 38.23 30.38 33.72 26.49 35.01
Qwen2-VL-2b [50] 11 24.60 19.43 38.03 40.63 18.84 14.09 7.81 7.07 17.82 11.14 27.55 26.21 25.25 31.20 28.22 54.12 49.11 21.75 25.27 12.47 23.92 27.62 24.83 14.85
Qwen2-VL-7b [50] 6 30.74 27.52 35.97 35.22 20.83 12.88 28.68 29.95 28.21 28.45 20.44 35.37 20.25 5.69 37.03 59.71 52.38 30.25 38.46 41.00 22.04 28.49 22.52 38.38
Qwen2.5-VL-7b [6] 2 33.07 28.75 31.33 33.66 21.99 14.97 42.88 37.73 23.83 23.64 22.97 33.33 28.75 6.83 40.27 58.24 51.49 44.75 50.00 32.13 33.87 32.85 27.15 31.93
LLaVA-v1.5-7b [36] 12 23.65 10.85 5.17 12.53 17.37 11.34 7.25 5.26 18.73 9.12 26.50 24.43 26.75 28.31 34.09 51.18 52.38 34.25 24.18 26.87 34.68 29.94 22.52 30.81
LLaVA-v1.6-7b [37] 13 13.21 8.53 12.14 0.00 20.35 0.27 10.76 0.41 24.27 0.00 4.79 6.62 7.75 0.00 20.18 51.76 7.74 6.25 32.14 6.37 39.52 10.47 21.52 5.88
InternVL2.5-8B+SPAR-mix - 63.25 65.53 81.53 79.22 38.25 35.51 78.93 79.18 68.13 63.50 63.01 78.88 73.00 37.14 61.29 86.47 79.76 64.00 69.23 59.00 47.31 50.00 42.38 53.50
Table 4: Performance of different models on SPAR-Bench. Shaded cells indicate the best scores in each category. The highest, second-highest, and third-highest scores in each category are highlighted with light red, light orange, and light yellow, respectively. SPAR-Bench (tiny) refers to a subset of the full benchmark, where 50 questions are sampled per task.
4.2 Evaluation on SPAR-Bench
Our model performs well on out-of-domain spatial benchmarks. We further evaluate it on our SPAR-Bench and assess the spatial capabilities of other large language models.
Human vs. model performance.
Tab. 4 shows that human performance far exceeds all models across all difficulty levels. The average human score is 67.27, with 55.31 on low-level tasks, 72.32 on middle-level tasks, and 76.22 on high-level tasks. Among baseline models (excluding our fine-tuned version), Qwen2.5-VL-72B performs best, but still falls far below human level, with an average score of 39.40 (35.35 on low-level, 23.05 on middle-level, and 48.44 on high-level tasks). This highlights that even the strongest models struggle with complex spatial reasoning, particularly as task difficulty increases.
Limitations of existing models.
The overall performance trend indicates that current models lack strong spatial perception and generalization ability for higher-order reasoning. The significant drop in scores from low-level to middle-level tasks suggests that models struggle to transition from basic spatial perception to more complex spatial relationships. For example, GPT-4o drops from 29.25 (low-level) to 24.93 (middle-level), and Qwen2.5-VL-72B drops from 35.35 to 23.05, highlighting difficulties in handling increasing spatial complexity. Additionally, performance in high-level tasks remains low, with most models scoring below 50, showing that multi-view and global spatial reasoning remain major challenges.
Impact of data augmentation.
Fine-tuning InternVL2.5-8B on our dataset significantly improves performance across all levels. The fine-tuned model scores 65.53 on low-level, 63.01 on middle-level, and 61.29 on high-level tasks, showing notable gains in spatial reasoning. For example, Depth-OC increases from 25.78 (InternVL2.5-8B) to 81.53, and View Change Inference improves from 6.32 to 37.14. These improvements suggest that high-quality spatial data enhances model performance, particularly in multi-view reasoning. However, the persistent gap between models and human performance indicates that data augmentation alone is insufficient, and further advances in multi-view fusion and spatial representation learning are needed.
SQA3D [42] ScanQA [4]
Methods EM@1 BLEU-4 CiDEr
3D-LLM [25] - 12.0 69.4
Chat-3D v2 [27] 54.7 14.0 87.6
LEO [29] 50.0 13.2 101.4
LL3DA [14] - 13.5 76.8
Scene-LLM [23] 54.2 12.0 80.0
SPAR-mix 58.1 15.3 90.7
Table 5: Comparison of SQA3D and ScanQA performance across different LLM-based models.
Methods Acc@0.25 Acc@0.5
ScanRefer  [11] 37.3 24.3
MVT [30] 40.8 33.3
ViL3DRel [13] 47.9 37.7
3D-LLM [25] 30.3 -
Chat-3D v2 [27] 35.9 30.4
Grounded 3D-LLM [15] 47.9 44.1
SPAR-mix 48.8(31.9) 43.1(12.4)
Table 6: Comparison of ScanRefer performance across different models. The content in “()” indicates results without refine.
4.3 Evaluation on 3D-specific tasks
In addition to the previously discussed spatial benchmarks, there are also tasks specifically defined in 3D scenes, which typically involve questions related to the entire scene.
After pretraining the VLM using our dataset, we perform supervised fine-tuning (SFT) on 3D tasks to further investigate the spatial capabilities of the VLM.
Spatial QA.
We use two benchmarks: ScanQA, which focuses on object attributes and spatial relationships within a scene, and SQA3D, a spatial benchmark where the task is to answer specific questions based on a given viewpoint description within the scene. This requires the model to demonstrate spatial imagination. We treat this as a multi-view task during training. As shown in Tab. 5, despite not incorporating 3D information into the model, it still achieves the highest score when compared to the 3D LLMs.
3D grouding.
We evaluate the 3D grounding capability of our model using the ScanRefer dataset. ScanRefer requires the model to locate objects in a scene based on textual descriptions. As shown in the Tab. 6, the values in parentheses represent the performance without proposal refinement. We observed that the VLM’s depth predictions may sometimes be inaccurate when the distance is large due to the absence of 3D information. However, after refining, the results perform well. These results underline the potential of leveraging a VLM for grounding, showcasing its ability to perform well without the need for 3D information, while also tackling a more complex generative task.
CV-Bench [49] SPAR-Bench
ViT SPAR Ratio 2D 3D Low Medium High
Frozen 60% 68.68 89.67 66.87 64.28 55.43
Unfrozen 60% 73.00 91.10 67.01 44.04 51.86
Unfrozen 40% 74.13 89.09 66.97 42.21 48.19
Table 7: Performance comparison of different ViT configurations. Models are trained with varying SPAR ratios and evaluated on CV-Bench (2D, 3D) and SPAR-Bench.
4.4 Ablation study
We evaluate the impact of ViT training strategies and SPAR data proportions on InternVL2.5-4B, trained with 2M samples from SPAR (3D) and EMOVA (2D). Tab. 7 presents results on CV-Bench (2D, 3D) and SPAR-Bench (Low, Middle, High levels).
Effect of freezing vs. unfreezing ViT.
Comparing the first two rows with 60% SPAR, unfreezing ViT improves CV-Bench 2D (68.68 to 73.00) and 3D (89.67 to 91.10), but significantly degrades SPAR-Bench Middle (64.28 to 44.04) and High (55.43 to 51.86). This suggests that fine-tuning ViT benefits general vision tasks but weakens spatial reasoning, particularly in multi-view scenarios.
Effect of SPAR-7M ratio.
Comparing the second and third rows with Unfrozen ViT, reducing the SPAR ratio from 60% to 40% slightly improves CV-Bench 2D (73.00 to 74.13) but reduces CV-Bench 3D (91.10 to 89.09). SPAR-Bench performance further drops, particularly in middle level (44.04 to 42.21) and high level (51.86 to 48.19) tasks, indicating that higher 3D data proportions enhance multi-view spatial reasoning.
Findings.
Freezing ViT preserves spatial understanding, while unfreezing benefits 2D tasks but harms multi-view reasoning. Higher SPAR ratios improve 3D reasoning, while increasing 2D data slightly benefits 2D vision at the cost of 3D perception.
5 Conclusion
We introduce SPAR-7M, a large-scale dataset constructed from thousands of scenes, specifically designed to enhance spatial reasoning. To provide a rigorous evaluation framework, we propose SPAR-Bench, a benchmark that offers a more comprehensive assessment of spatial understanding, supporting both single-view and multi-view inputs. Experimental results demonstrate that pretraining on SPAR-7M alongside general datasets enables models to achieve SOTA performance on VSI-Bench and CV-Bench 3D. Furthermore, fine-tuning on 3D task-specific datasets yields competitive results, highlighting the effectiveness of our dataset in bridging the gap between 2D and 3D spatial reasoning.