### v4

# From Flatland to Space: Teaching Vision-Language Models to Perceive and Reason in 3D

Abstract
Recent advances in LVLMs have improved vision-language understanding, but they still struggle with spatial perception, limiting their ability to reason about complex 3D scenes. Unlike previous approaches that incorporate 3D representations into models to improve spatial understanding, we aim to unlock the potential of VLMs by leveraging spatially relevant image data. To this end, we introduce a novel 2D spatial data generation and annotation pipeline built upon scene data with 3D ground-truth. This pipeline enables the creation of a diverse set of spatial tasks, ranging from basic perception tasks to more complex reasoning tasks. Leveraging this pipeline, we construct SPAR-7M, a large-scale dataset generated from thousands of scenes across multiple public datasets. In addition, we introduce SPAR-Bench, a benchmark designed to offer a more comprehensive evaluation of spatial capabilities compared to existing spatial benchmarks, supporting both single-view and multi-view inputs. Training on both SPAR-7M and large-scale 2D datasets enables our models to achieve state-of-the-art performance on 2D spatial benchmarks. Further fine-tuning on 3D task-specific datasets yields competitive results, underscoring the effectiveness of our dataset in enhancing spatial reasoning.
Figure 1: Overview of our Spatial Perception And Reasoning (SPAR) dataset and benchmark. Our dataset is sourced from 4,500 scenes and comprises 33 spatial tasks spanning single-view, multi-view, and video settings. Our benchmark includes over 7,000 carefully curated high-quality samples to comprehensively evaluate the spatial understanding capabilities of existing models.
1 Introduction
Recent advances in Large Vision-Language Models (LVLMs) have significantly enhanced vision-language understanding [1, 2, 3, 4, 5]. However, to empower VLMs to better interact with the physical world, spatial understanding is essential, serving as a cornerstone of embodied intelligence. Nevertheless, current VLMs exhibit limited spatial perception, often struggling with comprehensive understanding and reasoning about spatial relationships in complex scenes [6, 7, 8, 2, 9, 10].
A recent line of research explores enhancing spatial understanding by incorporating 3D representations of scenes [8, 9, 10, 11, 12]. However, these methods encounter several challenges. First, 3D data are scarce, with suboptimal quality and uneven distribution, making them unsuitable for large-scale pretraining. Second, designing a model (e.g., a point cloud encoder) to effectively capture 3D information is inherently challenging, particularly since complete 3D representations are often inaccessible in real-world scenarios. Moreover, these approaches modify the network architecture, hindering seamless integration with existing VLM frameworks [5, 3].
This raises the question: Is it truly necessary to incorporate 3D representations into VLMs? For humans, spatial understanding of the relationship between surrounding objects is innate, relying on the ability to observe and implicitly reconstruct the entire space from 2D observations. Motivated by this capability, we seek to empower VLMs to acquire spatial understanding directly from large-scale 2D data (e.g., image inputs), enabling them to perform spatial reasoning without explicit 3D supervision. To facilitate this, we construct a spatially relevant dataset tailored for training and evaluation. Additionally, to enable VLMs to tackle a broader range of 3D tasks, we design a multi-view-based 3D grounding approach.
We categorize spatial understanding into two levels: Spatial Perception, which encompasses fundamental spatial concepts such as depth, distance, relative positions, and camera viewpoints, representing low-level cognitive abilities; and Spatial Reasoning, which involves higher-order spatial understanding, requiring the ability to infer and construct 3D scenes implicitly.
Building on these two levels of spatial understanding, we introduce SPAR-7M, a dataset comprising 33 tasks that span a broad range of complexities. To facilitate systematic learning and evaluation, we organize all tasks into three levels—low, medium, and high—reflecting a natural progression from spatial perception to spatial reasoning. While previous datasets have predominantly focused on semantic descriptions, SPAR-7M places equal emphasis on numerical representations, enabling models to develop a more accurate and grounded understanding of 3D space. For example, we include fundamental perception tasks such as depth and distance estimation. We argue that under multi-view inputs, pixel- or feature-level matching is a foundational capability that supports higher-level spatial reasoning. Accordingly, many of our tasks—such as cross-view object matching and viewpoint transformation prediction—are explicitly designed to evaluate and strengthen this essential skill.
To generate the necessary training data, we introduce a 2D data generation and annotation pipeline that derives rich supervision from 3D scene data. The resulting dataset, SPAR-7M, provides precise 3D ground-truth annotations, significantly improving both task diversity and reliability. Unlike existing approaches that rely solely on 2D annotations [6, 7], our dataset enables a more structured and accurate assessment of spatial understanding.  Table 1 illustrates the diversity of input formats and task types within our dataset.
Dataset Perception Reasoning Single view Multi-view Video Task types QA quantity
ScanQA [13] - 27k
SQA3D [14] 16 33k
3D-LLM [8] 9 659k
LEO [15] 9 513k
SceneVerse [16] 3 2.5M
MMScan [17] 20 6.9M
SPAR-7M 33 7M
Table 1: Spatial understanding datasets comparison. Multi-view uses 3–5 images per QA pair.
Despite the growing body of spatial benchmarks, many focus predominantly on high-level spatial reasoning, often overlooking fundamental spatial perception [18, 19, 17]. Furthermore, some benchmarks, particularly those using internet-collected images, struggle to effectively evaluate multi-view tasks, while others based on video inputs are not well-aligned with typical model inputs [20]. To address these limitations, we select 20 tasks from the SPAR-7M validation set. After carefully manual validation and filtering, we created SPAR-Bench, a comprehensive benchmark consisting of 7,207 questions that span tasks from basic perception to complex reasoning. SPAR-Bench supports both single- and multi-view inputs, offering a more holistic and diverse evaluation of spatial understanding.
Our main contributions are as follows: (i). We present a 2D spatial QA generation pipeline using 3D ground truth to create both single- and multi-view data. This yields tasks covering numerical and semantic reasoning, forming a more comprehensive spatial understanding dataset. (ii). We present SPAR-Bench, a benchmark for evaluating the spatial capabilities of existing models. (iii). By training models on both SPAR-7M and large-scale 2D internet datasets, we achieve state-of-the-art performance across multiple 2D spatial benchmarks. Moreover, even with 2D-only inputs, further fine-tuning on 3D task-specific datasets yields competitive results.
2 Related works
Datasets for spatial understanding.
Existing multimodal datasets primarily provide single-view or limited multi-view images, often designed for image-text retrieval or basic visual question answering, making them insufficient for advanced 3D reasoning [21, 22, 23, 24]. Some datasets attempt to address this by collecting large-scale images from the internet, but they lack precise 3D annotations, relying instead on estimation models, which limits their effectiveness in multi-view and spatial reasoning tasks [6, 7]. Others incorporate image sequences or 3D scene representations for object counting, spatial relationship reasoning, or grounding, but they mainly focus on high-level semantics while neglecting low-level geometric cues such as depth, distance, and camera parameters [13, 14, 25]. Recent efforts [26] explore spatial reasoning from a robotics perspective, emphasizing reference-frame understanding and grounded spatial relations in real-world scenes. While highly relevant to embodied agents, these datasets focus on a narrower set of reasoning types and lack the task diversity and cognitive stratification we aim for. To overcome these limitations, we propose a 3D-ground-truth-based data generation pipeline capable of rendering single-view, multi-view, and video-based data. Our approach ensures precise spatial supervision, covering a broad spectrum of tasks from basic perception to advanced reasoning, structured into three difficulty levels. This provides a more comprehensive and scalable solution compared to existing datasets.
Benchmarks for evaluating spatial understanding.
Existing spatial understanding benchmarks mainly focus on high-level reasoning while often neglecting fundamental spatial perception tasks [19, 18, 17, 13, 14, 20]. Most of them rely on either single-view images or video inputs, with limited consideration of multi-view settings, despite their ability to better represent local 3D structures. Single-view images restrict the complexity of the questions and thus provide an incomplete assessment of VLMs’ spatial reasoning capabilities. While some benchmarks use video inputs [20], many VLMs are not optimized for video processing, but this does not necessarily mean they lack spatial reasoning abilities. To address these limitations, we introduce SPAR-Bench, a more comprehensive benchmark that encompasses both fundamental and advanced spatial tasks. It supports both single- and multi-view inputs, enabling a more systematic and reliable evaluation of spatial understanding.
3 Method
Our research aims to equip VLMs with robust and accurate spatial understanding, focusing on both fundamental 3D perception and higher-order spatial reasoning and imagination. To address the challenges posed by limited data, we propose a novel data generation pipeline that creates a diverse range of spatial understanding tasks. Our pipeline leverages 3D scenes from ScanNet [27], ScanNet++ [28], and Structured3D [29], ensuring a broad and diverse representation of indoor environments. Along with this, we introduce a new benchmark to provide a comprehensive evaluation of models’ spatial capabilities. In the following sections, we first detail our data generation pipeline (section 3.1) and the key characteristics of SPAR-7M (section 3.2). We then introduce SPAR-Bench (section 3.3), followed by an analysis of data quality and the manual filtering process used to ensure benchmark reliability (section 3.4). Additionally, in section 3.5, we present a multi-view-based 3D grounding approach designed to enhance VLMs’ ability to tackle a broader range of 3D tasks.
3.1 3D-driven data generation pipeline
This section presents our data generation pipeline (detailed in fig. 2), which takes as input a scene mesh, 3D object bounding boxes, camera parameters, and image sequences. Through filtering, 3D information extraction, and task construction, we produce diverse spatial understanding data.
Figure 2: Data generation pipeline. Our pipeline consists of three main parts: (1) Filtering redundant images by sampling approach. (2) Extracting scene 3D information and creating a unified data structure to support multi-view cross-retrieval and annotation. (3) Leveraging metadata to automatically generate a variety of tasks and forms of QA data.
Filtering and subsampling.
In the first step, we focus on reducing data redundancy by filtering out images with minimal camera movement. Specifically, each image is associated with a camera pose. If an image’s camera position is within a distance threshold of a retained image and its orientation differs by less than an angular threshold, it is considered redundant and removed. The whole process is similar to keyframing and the implementation details can be found in the appendix 6.2.
Extracting 3D information.
In the second step, we extract global 3D information from each scene and organize it into a unified data structure for cross-view retrieval and downstream annotation. This structure includes two types of entities: image records and object records, which together create a consistent mapping between image and object data in world coordinates. An image record stores image-level metadata, including index, camera parameters, resolution, and visible objects. Each object is linked to its 3D bounding box, 3D center in world coordinates, and 2D projection. An object record contains object-level metadata, including ID, category, and a list of associated image indices. This compact representation facilitates multi-dataset alignment and supports scalable task generation. Technical details are provided in appendix 6.3.
Multi-tasks generation.
In the third step, we use the complete 3D scene information to automatically generate tasks and corresponding QA pairs. Each task is configured with key details such as task type (e.g., distance prediction, spatial relations), input image count, question format, and the object types involved (object-to-object or object-to-observer relations). For each task, we follow a specific pipeline to generate annotations, ensuring flexibility to extend or customize tasks as needed. After generating the annotations, placeholders in the template (e.g., [label_A], [label_B], [left_right]) are replaced with actual object labels and computed spatial information. The templates are initially created by humans, expanded with the help of GPT, and then refined through manual selection to ensure diversity and accuracy. Detailed information about task pipelines, QA templates, and GPT prompting can be found in the appendix 6.4.
Figure 3: Dataset and benchmark visualization. We present illustrative samples from SPAR-7M and SPAR-Bench to highlight the diversity of spatial tasks and input types. The examples have been simplified for clarity; Full task formats are provided in the released dataset and Appendix 6.5.
3.2 Spatial dataset construction
Based on our data generation pipeline, we construct SPAR-7M, a large-scale dataset designed for spatial understanding. SPAR-7M consists of 4,000+ indoor 3D scenes from diverse sources, covering 33 task types with over 7 million QA pairs. Each QA pair is paired with camera parameters, and depth, enabling both multi-view reasoning and 3D spatial inference. The dataset includes tasks ranging from basic spatial perception to high-level spatial reasoning (detailed in fig. 4).
To evaluate diverse aspects of spatial understanding, we design tasks with three input formats: single-view, multi-view, and video. Single-view tasks assess spatial inference from one image, while multi-view tasks require cross-view integration. Video inputs introduce scene-level reasoning and long-range spatial perception. This diversity enables a comprehensive evaluation across settings. Tasks are further categorized by spatial relation type: object-object (OO) and object-camera (OC). OO tasks focus on inter-object positioning (e.g., left/right/behind), while OC tasks involve viewpoint-dependent relationships. For multi-view inputs, one image is randomly chosen as the primary viewpoint for reasoning.
Figure 4: Dataset Statistics. Our Dataset can be categorized into high, medium, and low three levels, totaling 33 subtasks.
3.3 Spatial benchmark construction
SPAR-Bench is constructed to evaluate models across a wide range of spatial tasks, from low-level geometric estimation to high-level relational reasoning. It includes 20 tasks sampled from the SPAR-7M test set, each with 400 QA pairs and manually verified for quality, resulting in a total of 7207 QA samples. The tasks span both single-view and multi-view settings and are formulated as either numerical prediction or classification, evaluated using Mean Relative Absolute Error (MRA) or accuracy, respectively. While the full dataset includes 33 spatial tasks, we select 20 for evaluation. Tasks requiring video input are excluded to disentangle spatial understanding from temporal modeling. Our focus is on evaluating how well models reason about 3D structure, geometry, and viewpoint relations from one or more views, independent of motion or continuity cues that video inputs may introduce. This also avoids overlap with existing video-based spatial benchmarks, and positions our evaluation as a complementary perspective focused purely on spatial reasoning.
To cover different facets of spatial understanding, the benchmark includes tasks such as depth estimation, distance comparison, object relation reasoning, and spatial imagination. Single-view tasks assess inference from a static observation, while multi-view tasks require reasoning across perspectives. A full task list is available in appendix 6.1, with QA visualizations in appendix 6.5, and dataset examples shown in fig. 3.
3.4 Data quality check
To ensure the reliability of our benchmark, we conduct a human verification process. Specifically, we sample 400 questions per task in the validation set, forming an 8,000-question benchmark. These questions undergo manual inspection, taking approximately 140 human hours. After filtering out problematic cases, 7,207 questions were retained, resulting in a dataset acceptance rate of 90.1%.
During verification, we focus on identifying vague question stems, incorrect or misleading choices, ambiguous answers, and factual inconsistencies. Errors are marked, discussed, and either corrected or removed. This rigorous process ensures that our benchmark maintains high quality and reliability.
The 90.1% acceptance rate indicates that the vast majority of samples were logically valid, clearly phrased, and visually answerable. In contrast, the 67.27 average human score table 4 reflects the intrinsic difficulty of the proposed spatial reasoning tasks—particularly those involving occlusion, viewpoint ambiguity, or subtle geometric cues. Rather than suggesting flaws in annotation quality, this gap underscores the challenging nature of our benchmark and its ability to expose limitations in both human and model spatial understanding.
Figure 5: Our 3D grounding pipeline via VLM.
3.5 3D grounding with VLMs
Expressing an object’s spatial position within a scene is a crucial capability. Previous methods primarily relied on point cloud regression or classifying proposal bounding boxes. For video inputs, temproal grounding in VLM [30, 31] is widely explored, [32] use VLM as a frame selector to achieve 3D grounding with a series of tools. In contrast, we propose a novel 3D grounding approach that reformulates this task as a text generation problem, enabling seamless integration with VLMs. Specifically, we transform the 3D grounding task into a frame selection followed by mono-3D grounding. In the case of multi-view input, an object can be localized within a visible frame, allowing us to infer its UV coordinates, depth, and size, achieving accurate 3D localization through the camera pose of that frame. While this method already enables the localization of a 3D bounding box, we further introduce an optional refinement process. By optimally matching the predicted 3D bounding box with the proposal bounding box, we obtain the final refined result. The overall grounding pipeline is illustrated in fig. 5.
4 Experiments
In this section, we assess the effectiveness of our dataset and investigate the potential upper limits of spatial reasoning in VLMs from the following perspectives: (i). Performance on existing spatial benchmarks: we examine how pretraining on SPAR-7M improves model performance on image-based spatial benchmarks; (ii). Evaluation on our in-domain benchmark SPAR-Bench: we examine how pretraining on SPAR-7M improves model performance on image-based spatial benchmarks; (iii). Fine-tuning on 3D-specific tasks: To investigate the generalization of our approach, we perform supervised fine-tuning (SFT) on 3D-specific benchmarks, including spatial QA and 3D grounding tasks.
Furthermore, in section 4.4, we analyze the impact of different data compositions and training strategies on model performance, offering insights into optimal configurations for enhancing spatial reasoning capabilities.
Pretraining.
To train a VLM with strong spatial capabilities while preserving generalization, we use a mixed pretraining dataset SPAR-mix. 40% of the data is uniformly sampled from our dataset, SPAR-7M, while the remaining 60% is uniformly sampled from general 2D data, sourced from the EMOVA configuration. This general data includes tasks such as image-text QA, chart understanding, and OCR. We pretrain the model on a total of 2 million samples, using the base model InternVL2.5-8B. Pretraining was performed on 64 NPUs for 4 days, resulting in a total of about 6,100 NPU-hours.
4.1 Evaluation on existing 2D spatial benchmarks
CV-Bench [33].
The Cambrian Vision-Centric Benchmark (CV-Bench) is a 2,638-sample dataset for evaluating 2D spatial reasoning and object counting. It repurposes vision datasets with natural language questions and uses visual prompts (e.g., bounding boxes) to reduce ambiguity and ensure reliable evaluation.
VSI-Bench [20].
The Visual-Spatial Intelligence Benchmark (VSI-Bench) evaluates spatial reasoning from egocentric video data, containing 5,000+ QA pairs from 288 real-world videos sourced from ScanNet [27], ScanNet++ [28], and ARKitScenes [34]. It covers diverse indoor environments and leverages object-level annotations for question generation. VSI-Bench includes eight tasks across three categories: configurational reasoning (object counting, relative position, route planning), measurement estimation (object and room size, absolute distance), and spatiotemporal reasoning (appearance order in videos). All questions undergo manual review to ensure accuracy.
To evaluate the effectiveness of our dataset, we compare models trained on SPAR-mix (ours with EMOVA-2M) against those trained solely on EMOVA-2M. Since our dataset focuses specifically on spatial perception and reasoning, we incorporate a subset of QA data from EMOVA to maintain performance on general-purpose vision-language tasks. Note that the Base VLM has already been pretrained on extensive data with optimized training strategies, making it an unfair baseline for assessing the incremental gain. Therefore, we consider Base VLM + EMOVA-2M as the proper baseline. These results validate the effectiveness of our dataset design and indicate that spatial supervision can be introduced without compromising general QA performance. As shown in table 2, our model achieves state-of-the-art results on CV-Bench 3D and VSI-Bench, demonstrating significant improvements in spatial understanding. While there are minor drops on general benchmarks such as MMBench and TextVQA—mostly within 1%—this is expected, as our data emphasizes spatial reasoning rather than fine-grained vision-language alignment.
VSI-Bench [20] CV-Bench   [33] BLINK [18] 3DSRBench [19] Seed-Image [21] MME [35] MMBench [24] RealWorldQA [36] TextVQA [37]
Methods 2D 3D
GPT-4v [38] - 64.3 73.8 51.14 - 71.6 1927 75.0 61.4 77.4
GPT-4o [38] 34.0 - - 60.04 45.3 77.1 2310 83.4 75.4 -
Cambrain-8B [33] - 72.3 72.0 - - 74.7 1547 75.9 64.2 77.8
LLaVA-OV-7B [4] 25.3 - - 48.2 44.1 - 1998 80.8 66.3 -
InternVL2-8B [39] 34.6 - - - - 75.4 2215 81.7 - 77.4
Base VLM 32.4 74.20 78.50 46.61 58.33 76.53 2323 84.45 66.67 68.73
+EMOVA-2M 24.5 66.27 64.83 42.40 55.25 73.8 2186 80.24 63.14 63.78
+SPAR-mix 41.1 72.25 89.08 43.92 57.48 73.2 2163 79.90 64.71 62.91
Table 2: Comparison of model performance on various spatial understanding and 2D general benchmarks. The base VLM is InternVL2.5-8B, and we use dynamic patch of 3. Note: Our main comparison is between +EMOVA-2M and +SPAR-mix, as both build upon the same Base VLM and general QA data. For detailed discussion, see section 4.1.
QA source Hypersim [40] SUNRGBD [41] nuScenes [42]
Base VLM 73 87 62
+SPAR-mix 93 96 80
Table 3: Performance improvement on out-of-domain datasets in CV-Bench 3D.
Out-of-domain generalization ability.
We also find that pretraining with our dataset enhances the model’s spatial understanding across diverse, non-homogeneous datasets. Although our dataset is primarily derived from indoor scene datasets such as ScanNet [27], ScanNet++ [28], and Structured3D [29], the pretrained model shows consistent improvements across all tasks constructed with datasets different from ours in CV-Bench 3D, as shown in table 3.
Notably, our method achieves an 18-point gain on the outdoor nuScenes dataset, highlighting strong generalization beyond indoor scenes. This improvement stems from the spatial reasoning skills emphasized in our dataset—many CV-Bench 3D questions focus on relative distance comparisons, similar to tasks like DistI-OO and ObjRel. Since these tasks rely on 3D geometric cues rather than appearance or domain-specific semantics, the learned spatial priors transfer effectively across diverse environments.
Methods
Rank
Avg.
Low
Depth-OC
Depth-OC-MV
Depth-OO
Depth-OO-MV
Dist-OC
Dist-OC-MV
Dist-OO
Dist-OO-MV
Medium
PosMatch
CamMotion
ViewChgI
High
DistI-OO
DistI-OO-MV
ObjRel-OC-MV
ObjRel-OO
ObjRel-OO-MV
SpImag-OC
SpImag-OC-MV
SpImag-OO
SpImag-OO-MV
Baseline
Chance Level (Random) - - - - - - - - - - - - 22.65 24.50 - 25.09 23.82 22.02 31.25 25.27 22.16 25.81 24.42 24.17 26.89
Chance Level (Frequency) - 32.74 31.19 43.09 43.51 17.38 13.05 41.90 30.99 27.40 32.17 38.25 29.01 26.75 59.00 32.29 52.94 50.60 28.25 26.92 26.59 26.34 26.74 26.49 25.77
SPAR-Bench (tiny) API
Human Level 1 67.27 55.31 72.75 74.25 28.75 36.25 78.25 52.25 66.5 33.50 72.32 92 64 60.97 76.22 80 94 70 92 80 78 82 50 60
GPT-4o [38] 3 36.39 29.25 53.80 45.00 15.00 13.60 37.40 34.40 23.40 24.40 24.93 30 16 28.80 45.11 64 64 58 46 46 32 44 30 22
Claude-3.7-Sonnet [43] 5 21.77 25.43 41.00 45.40 11.20 12.20 42.60 19.60 26.00 5.40 7.33 16 6 0.00 23.33 40 48 22 36 14 12 20 6 12
Qwen2-VL-72B [44] 4 35.62 35.28 45.40 49.80 13.80 10.00 54.60 49.40 36.80 22.40 23.39 42 18 10.16 40.00 60 68 50 38 44 18 28 18 36
Qwen2.5-VL-72B [45] 2 39.40 35.35 53.20 46.80 17.80 29.00 49.60 57.40 14.40 14.60 23.05 40 16 13.16 48.44 74 74 60 56 50 20 34 24 44
SPAR-Bench (full)
InternVL2-2B [39] 10 28.06 21.74 18.06 24.81 23.20 20.97 19.47 19.95 26.83 20.61 22.83 39.69 23.00 5.81 35.42 51.18 55.95 46.00 31.59 23.82 36.02 34.30 17.55 22.41
InternVL2-4B [39] 4 32.01 28.94 23.94 27.22 20.00 18.12 42.57 40.16 31.29 28.18 29.16 49.87 21.00 16.62 35.70 56.76 55.36 40.25 36.81 25.21 28.76 32.27 21.19 24.65
InternVL2-8B [39] 3 33.02 26.83 25.75 30.88 20.67 20.78 39.03 36.19 19.15 22.19 36.49 63.36 28.00 18.11 37.37 64.71 54.46 42.75 37.36 26.32 34.14 31.10 20.86 24.65
InternVL2.5-2B [46] 8 30.14 25.79 39.67 39.72 12.12 15.03 30.94 29.59 20.22 19.02 22.93 37.91 24.25 6.64 36.41 51.47 56.85 50.25 33.79 24.10 27.15 35.17 26.49 22.41
InternVL2.5-4B [46] 7 30.55 25.66 29.06 32.97 21.77 16.83 20.84 26.85 28.13 28.79 29.75 47.07 33.25 8.92 35.16 54.12 58.93 35.50 29.67 34.63 24.73 31.39 19.21 28.29
InternVL2.5-8B [46] 1 36.28 29.46 25.78 29.31 23.79 18.76 46.82 42.68 22.62 25.89 31.88 61.32 28.00 6.32 43.80 59.71 56.85 51.75 44.23 41.55 36.56 41.57 22.52 39.50
LLaVA-OV-0.5B [4] 9 29.48 30.14 49.22 42.72 18.04 14.92 31.48 25.67 28.98 30.10 15.89 24.43 21.75 1.50 33.42 50.88 50.00 32.00 27.75 26.04 30.91 34.01 24.50 24.65
LLaVA-OV-7B [4] 5 31.20 21.79 30.33 26.94 18.58 13.87 10.43 13.64 31.24 29.29 26.13 38.68 30.25 9.47 40.14 56.47 55.06 37.25 48.63 38.23 30.38 33.72 26.49 35.01
Qwen2-VL-2b [44] 11 24.60 19.43 38.03 40.63 18.84 14.09 7.81 7.07 17.82 11.14 27.55 26.21 25.25 31.20 28.22 54.12 49.11 21.75 25.27 12.47 23.92 27.62 24.83 14.85
Qwen2-VL-7b [44] 6 30.74 27.52 35.97 35.22 20.83 12.88 28.68 29.95 28.21 28.45 20.44 35.37 20.25 5.69 37.03 59.71 52.38 30.25 38.46 41.00 22.04 28.49 22.52 38.38
Qwen2.5-VL-7b [45] 2 33.07 28.75 31.33 33.66 21.99 14.97 42.88 37.73 23.83 23.64 22.97 33.33 28.75 6.83 40.27 58.24 51.49 44.75 50.00 32.13 33.87 32.85 27.15 31.93
LLaVA-v1.5-7b [47] 12 23.65 10.85 5.17 12.53 17.37 11.34 7.25 5.26 18.73 9.12 26.50 24.43 26.75 28.31 34.09 51.18 52.38 34.25 24.18 26.87 34.68 29.94 22.52 30.81
LLaVA-v1.6-7b [48] 13 13.21 8.53 12.14 0.00 20.35 0.27 10.76 0.41 24.27 0.00 4.79 6.62 7.75 0.00 20.18 51.76 7.74 6.25 32.14 6.37 39.52 10.47 21.52 5.88
InternVL2.5-8B+SPAR-mix - 63.25 65.53 81.53 79.22 38.25 35.51 78.93 79.18 68.13 63.50 63.01 78.88 73.00 37.14 61.29 86.47 79.76 64.00 69.23 59.00 47.31 50.00 42.38 53.50
Table 4: Performance of different models on SPAR-Bench. Shaded cells indicate the best scores in each category. The highest, second-highest, and third-highest scores in each category are highlighted with light red, light orange, and light yellow, respectively. SPAR-Bench (tiny) refers to a subset of the full benchmark, where 50 questions are sampled per task.
4.2 Evaluation on SPAR-Bench
Our model performs well on out-of-domain spatial benchmarks. We further evaluate it on our SPAR-Bench and assess the spatial capabilities of other large language models.
Human vs. model performance.
Table 4 shows that human performance consistently exceeds all models across difficulty levels, with a clear margin. While the strongest baseline (Qwen2.5-VL-72B) achieves competitive results, it still falls far short of human-level performance—particularly in tasks requiring spatial reasoning across views or precise geometric estimation.
While performance across levels is not directly comparable due to different metrics (MRA vs. accuracy), comparing to human results reveals a clear trend: models perform moderately on low-level tasks (e.g., 35.4% for Qwen2.5-VL-72B), but drop significantly on middle-level tasks (23.1%), indicating difficulty in structured spatial reasoning. Performance partially rebounds on high-level tasks (48.4%), likely aided by language priors or pattern-based inference. Tasks involving viewpoint changes, spatial imagination, or indirect metric reasoning remain the most challenging, underscoring current limits in spatial generalization and multi-view consistency.
Impact of data augmentation.
Fine-tuning InternVL2.5-8B on SPAR-mix significantly boosts performance across all difficulty levels, reaching 65.53 (low), 63.01 (mid), and 61.29 (high). The largest gains occur in depth estimation and multi-view reasoning—e.g., View Change Inference improves from 6.32 to 37.14—demonstrating the effectiveness of structured spatial data in aligning geometric cues across viewpoints. Notably, the fine-tuned model surpasses human performance in depth prediction, and even the strongest pre-finetuning model exceeds humans on distance estimation. This is understandable, as these metrics are MRA-based and favor numerically optimized models over human intuition.
These results confirm that our dataset introduces valuable inductive bias for spatial reasoning. However, gains remain uneven, and a clear gap persists compared to human-level performance—highlighting the need for advances in multi-view fusion, spatial representation learning, and compositional reasoning.
SQA3D [14] ScanQA [13]
Methods 3D EM@1 BLEU-4 CiDEr
3D-LLM [8] - 12.0 69.4
Chat-3D v2 [49] 54.7 14.0 87.6
LEO [15] 50.0 13.2 101.4
LL3DA [9] - 13.5 76.8
Scene-LLM [12] 54.2 12.0 80.0
LLaVA-3D [10] 55.6 14.5 91.7
Video-3D LLM [50] 58.6 16.2 102.1
SPAR-mix 58.1 15.3 90.7
Table 5: Comparison of SQA3D and ScanQA performance. “3D” indicates whether the model is infused with 3D information.
Methods Acc@0.25 Acc@0.5
ScanRefer [51] 37.3 24.3
MVT [52] 40.8 33.3
ViL3DRel [53] 47.9 37.7
3D-LLM [8] 30.3 -
Chat-3D v2 [49] 35.9 30.4
Grounded 3D-LLM [54] 47.9 44.1
LLaVA-3D [10] 54.1 42.4
Video-3D LLM [50] 58.1 51.7
SPAR-mix 48.8 (31.9) 43.1 (12.4)
Table 6: Comparison of ScanRefer performance across different models. The content in “()” indicates results without refine.
4.3 Evaluation on 3D-specific tasks
In addition to the previously discussed spatial benchmarks, there are also tasks specifically defined in 3D scenes, which typically involve questions related to the entire scene. After pretraining the VLM using our dataset, we perform supervised fine-tuning (SFT) on 3D tasks to further investigate the spatial capabilities of the VLM.
Spatial QA.
We use two benchmarks: ScanQA, which focuses on object attributes and spatial relationships within a scene, and SQA3D, a spatial benchmark where the task is to answer specific questions based on a given viewpoint description within the scene. This requires the model to demonstrate spatial imagination. We treat this as a multi-view task during training. As shown in table 6, despite not incorporating 3D information into the model, it still achieves the competitive scores when compared to the 3D-enhanced LLMs.
3D grouding.
We evaluate the 3D grounding capability of our model using the ScanRefer dataset. ScanRefer requires the model to locate objects in a scene based on textual descriptions. As shown in the table 6, the values in parentheses represent the performance without proposal refinement. We observed that the VLM’s depth predictions may sometimes be inaccurate when the distance is large due to the absence of 3D information. However, after refining, the results perform well.
These results underline the potential of leveraging a VLM for grounding, showcasing its ability to perform well without the need for 3D information, while also tackling a more complex generative task.
CV-Bench [33] SPAR-Bench
SPAR Ratio 2D 3D Low Medium High
40% 74.13 89.09 66.97 42.21 48.19
60% 73.00 91.10 67.01 44.04 51.86
Table 7: Performance comparison of different SPAR ratios.
4.4 Ablation study
We evaluate the impact of SPAR data proportions on InternVL2.5-4B, trained with 2M samples from SPAR (3D) and EMOVA (2D), using an unfrozen ViT. We adopt ViT fine-tuning as it improves overall performance and better adapts to the diverse spatial structures in SPAR-7M. Table 7 presents results on CV-Bench (2D, 3D) and SPAR-Bench (Low, Middle, High levels).
Effect of SPAR ratio.
With an unfrozen ViT, reducing the SPAR-7M ratio from 60% to 40% leads to a slight improvement on CV-Bench 2D (73.00 → 74.13), but results in performance drops on CV-Bench 3D (91.10 → 89.09) and across all levels of SPAR-Bench—particularly on the middle (44.04 → 42.21) and high (51.86 → 48.19) tasks. This indicates while 2D benchmarks may benefit from increased exposure to EMOVA-style data, an excessive reduction in the SPAR-7M proportion can cause the model to forget generalizable spatial priors. In contrast, complex spatial reasoning—especially under multi-view or relational conditions—relies more heavily on training signals tailored for spatial understanding. Maintaining or increasing the proportion of SPAR-7M is thus critical for reinforcing geometric alignment and ensuring spatial consistency across diverse viewpoints.
5 Conclusion
We introduce SPAR-7M, a large-scale dataset constructed from thousands of scenes, specifically designed to enhance spatial reasoning. To provide a rigorous evaluation framework, we propose SPAR-Bench, a benchmark that offers a more comprehensive assessment of spatial understanding, supporting both single-view and multi-view inputs. Experimental results show that pretraining on SPAR-7M alongside general datasets enables models to achieve state-of-the-art performance on VSI-Bench and CV-Bench 3D. Furthermore, fine-tuning on 3D task-specific datasets yields competitive results, highlighting the effectiveness of our dataset in bridging the gap between 2D and 3D spatial reasoning. While our tasks span diverse spatial challenges, they are primarily static QA-based and do not yet capture continuous spatial reasoning required in embodied or interactive settings.